Find the equation of the line through (2,a) and (3,b).
Answer by stanbon(75887)

Find the equation of the line through (2,a) and (3,b).
----
slope = (b-a)/(3-2) = b-a
-----------------------
Form: y = mx + k
Solve for "k":
a = (b-a)2 + k
a = 2b-2a + k
k = 3a-2b
----
Equation:
y = (b-a)x + (3a-2b)
==========================
Cheers,
Stan H.
=====================